If you intend to make $175,000 per year in the latter part of your career and you wish to maintain that in retirement, how much must you have saved at retirement? You expect to earn an average of 6% in retirement. You plan to retire at the age of 65 and you feel you need to conservatively maintain this income until the age of 100 years old. Total amount saved
Answer #1

Answer:,

Let us assume on retirement, annual withdrawal will be made at the beginning of year.

Annual withdrawal =$175,000

Number of years (Retirement duration) = 100 - 65 = 35 years

Rate of earnings during retirement = 6%

To get total retirement corpus (total amount saved) at the start of retirement, we will use PV function of excel:

= PV (rate,nper, pmt, fv, type)

= PV (6%, 35, -175000, 0, 1)

= $2,689,424.70

Total amount Saved (on retirement date) = $2,689,424.70